### v1

# Lightening the Load: A Cluster-Based Framework for A Lower-Overhead, Provable Website Fingerprinting Defense

Abstract
Website fingerprinting (WF) attacks remain a significant threat to encrypted traffic, prompting the development of a wide range of defenses. Among these, two prominent classes are regularization-based defenses, which shape traffic using fixed padding rules, and supersequence-based approaches, which conceal traces among predefined patterns. In this work, we present a unified framework for designing an adaptive WF defense that combines the effectiveness of regularization with the provable security of supersequence-style grouping. The scheme first extracts behavioural patterns from traces and clusters them into $(k,l)$-diverse anonymity sets; an early-time-series classifier (adapted from ECDIRE) then switches from a conservative global set of regularization parameters to the lighter, set-specific parameters. We instantiate the design as Adaptive Tamaraw, a variant of Tamaraw that assigns padding parameters on a per-cluster basis while retaining its original information-theoretic guarantee. Comprehensive experiments on public real-world datasets confirm the benefits. By tuning $k$, operators can trade privacy for efficiency: in its high-privacy mode Adaptive Tamaraw pushes the bound on any attacker’s accuracy below 30%, whereas in efficiency-centred settings it cuts total overhead by 99% compared with classic Tamaraw.
I Introduction
Tor is the leading low-latency anonymity network, relied upon by millions worldwide to shield their online activities from surveillance and censorship [1]. Despite its robust encryption and onion routing design, Tor leaks metadata such as packet sizes, timing, and directional patterns that adversaries can exploit. Over the years, website fingerprinting (WF) attacks have demonstrated that by analyzing these residual traffic features, even passive attackers can infer with high accuracy which webpages users are visiting [2, 3, 4]. Modern deep learning techniques employing transformer models [5] and multi-channel representations can now extract fine-grained patterns from traffic data, achieving higher recall and precision  [6, 7, 8, 9].
In response to these threats, a variety of defenses have emerged to mitigate WF attacks. One broad class of approaches is regularization-based defenses, which aim to conceal traffic patterns by enforcing packet transmission rules to reduce entropy. A common rule is a fixed, constant packet rate, achieved with delays and dummy packets [10, 11, 12]. These strategies often apply the same padding schedule to all sites, which leads to excessive overhead, particularly for bursty traffic. A notable example is Tamaraw [10], one of the few defenses to provide information-theoretic guarantees on attacker success. Tamaraw has not been broken; almost all defenses that lack such guarantees have been defeated by more powerful classifiers [13, 14, 15].
Another family of defenses aims to construct super-traces, forcing instances of similar webpages to produce the same pattern and thus hinder an attacker from inferring the correct webpage [2, 16, 17, 8]. However, these defenses are calibrated on a fixed set of sites. During an offline stage they partition that set into anonymity sets (clusters) with similar traffic characteristics and derive a canonical super‑trace for each cluster. At run time every page load is forced to follow the super-trace assigned to its destination page. A page that does not exist in the offline stage produces undefined behavior, but in practice, ordinary browsing covers far more destinations than any feasible reference set can capture.
In this work, we propose a hybrid website fingerprinting defense with the strengths of both regularization and supersequence approaches. Traffic loading under our defense begins with regularization: a global rate that protects early traffic without requiring prior knowledge of the destination. As the trace evolves, we switch to supersequence: a streaming early time series classifier assigns the live trace to a preconstructed anonymity set, after which the defense transitions to a lightweight regularization rate specific to that set. These sets are generated offline through clustering, subject to k-anonymity [18] and l-diversity [19]. Our construction satisfies an information theoretic upper bound on any adversary’s success.
Using this design framework, we create Adaptive Tamaraw, a novel extension of the original Tamaraw defense that retains its provable security while reducing overhead. We provide a rigorous analysis of Adaptive Tamaraw that bounds any attacking classifier’s maximum accuracy by the size and diversity of the anonymity set. We also evaluate our defense empirically: we test state-of-the-art website fingerprinting attacks to verify that the actual attack accuracy remains within our theoretical bounds while yielding lower overhead compared to the fixed-rate Tamaraw.
We summarize the main contributions of our work as follows:
  • We propose a general design framework for constructing provable website fingerprinting defenses that combines regularization defenses with dynamic clustering to adaptively adjust defense parameters in real time. As a concrete instantiation, we develop Adaptive Tamaraw, an extension of the original Tamaraw defense that preserves its information-theoretic guarantees while reducing bandwidth and latency overhead.
  • We provide a formal analysis that quantifies the privacy guarantees of Adaptive Tamaraw. Specifically, we derive upper bounds on the maximum achievable attack accuracy, independent of the underlying classifier, based on the size and diversity of each anonymity set.
  • Our experiments show that Adaptive Tamaraw offers flexible tunability, enabling practitioners to adjust the defense parameters to meet specific privacy and efficiency requirements. When configured for strong privacy, the defense reduces attack accuracy to well below 30%. In efficiency-focused settings, it achieves substantial reductions in the total overhead, up to 99% relative to the original Tamaraw. The implementation is publicly available at https://anonymous.4open.science/r/Adaptive-Tamaraw-F864.
II Background and Related Work
II-A Website Fingerprinting Attacks
Website fingerprinting attacks enable an on-path adversary to infer which website a user visits by analyzing observable traffic features such as packet sizes, timings, and ordering, even when the content is encrypted. Early methods relied on manually engineered features combined with classical classifiers like SVMs [20], k‑Nearest Neighbors [2], and k‑fingerprinting [4], the last of which uses random decision forests on manually extracted features.
[…]
II-B Website Fingerprinting Defenses
When designing defenses against website fingerprinting attacks, the literature has broadly followed two lines of work. The first focuses on empirical defenses that are evaluated primarily through experiments and often rely on heuristics or traffic obfuscation strategies to reduce attack success. The second aims to provide formal, information-theoretic guarantees that rigorously bound the adversary’s ability to classify webpages. Table I offers a comprehensive summary of both approaches, highlighting their key properties along with their strongest known attack accuracies. We now examine these two categories in more detail, beginning with empirical defenses.
II-B1 Empirical WF Defenses
A large body of website fingerprinting defenses rely on empirical evaluation to demonstrate effectiveness against specific attacks. These defenses often succeed in reducing fingerprinting accuracy and can be tuned for lower overhead. However, they typically lack formal guarantees and remain vulnerable to stronger or more adaptive adversaries. Empirical defenses can be broadly categorized into two main families:
Obfuscation-Based Defenses: These approaches inject randomness to obscure traffic patterns. WTF-PAD [13] adds dummy packets during idle periods, FRONT [14] perturbs early bursts with randomized padding, and Surakav [15] employs a GAN to simulate realistic timing. While effective against some attacks, they lack formal guarantees and are vulnerable to modern methods like Laserbeak [9] and RF [8]. As shown in Table I, these defenses still result in relatively high attack accuracies.
Regularization-Based Defenses: These defenses reduce leakage by shaping traffic into fixed or structured patterns. BuFLO [21] and CS-BuFLO [11] enforce constant rates, providing strong obfuscation but at high cost. RegulaTor [12] improves efficiency by reshaping only sensitive trace segments. However, most lack formal guarantees and thus can still leak expressive features.
While empirical defenses can reduce current attack success, their lack of formal guarantees limits robustness against future threats. In contrast, as we will discuss, Adaptive Tamaraw offers provable security (Section VI-D) and adaptively lowers overhead while preserving strong privacy guarantees.
TABLE I: Strongest published attacks against representative WF defences in the closed-world setting.
  • Defences without a provable formal bound are consistently broken by adversarial training, with the attacker’s accuracy remaining well above the 50 % threshold commonly associated with robust resistance [22].
  • Super-sequence defences achieve low attack accuracy but protect only those websites present in their training corpus, limiting their applicability in open-world browsing.
Category Defence Strongest published attack  Accuracy (%) Formal bound? Limited to Dataset?
Obfuscation WTF-PAD RF [8] 96.6 No No
FRONT LASERBEAK [9] 95.9 No No
Surakav LASERBEAK [9] 81.5 No No
Regularization RegulaTor RF [8] 67.4 No No
Tamaraw LASERBEAK [9] 25.3 Yes No
Walkie-Talkie RF [8] 93.9 Yes Yes
Super-sequence Super-Sequence Tik-Tok [7, 8] 29.18 Yes Yes
Palette RF [8, 22] 36.43 Yes Yes
 “Formal bound” indicates that the defence supplies an explicit, theoretical upper bound on an attacker’s success for any strategy. “Limited to Dataset” means the defence only applies to the websites contained in its training set, which is impractical for real-world browsing.
 Palette does not provide a closed-form mathematical bound; its “Yes” entry denotes that its cluster-anonymisation method yields structured leakage reductions that approach provable guarantees [22].
 Where two references appear, the first paper introduces the attack and the second reports its performance against the listed defence.
II-B2 WF Defenses With Formal Bounds
A number of defenses have sought not only to empirically reduce the success of website fingerprinting attacks but also to provide formal, provable bounds on security. Among these, Tamaraw [10] stands out as the first defense to derive an explicit upper bound on an adversary’s success probability. By construction, its uniform-rate padding guarantees that no attack, regardless of strategy, can exceed the bound as traces for different websites have the same rate and often have the same length. However, this theoretical rigor comes at a steep cost, and Tamaraw’s fixed-rate design introduces substantial overhead.
Other formal defenses attempt to balance rigor with efficiency. Walkie-Talkie [17], for example, reduces traffic uniqueness by pairing sites and enforcing collisions in their trace representations. While it provides a bounded adversarial success rate in closed-world settings, it assumes prior knowledge of the entire trace; an alternative randomized version avoids this assumption, but loses its adversarial bound.
Wang et al. introduced the Supersequence defense [2], which groups webpages into anonymity sets and pads each trace to match a common super-sequence, ensuring that all traces within a cluster are indistinguishable. Similarly, Palette [22] applies traffic cluster anonymization in real time by shaping traffic to match canonical burst patterns within each cluster, though it does not provide formal bounds. However, these supersequence-based methods are limited to the websites present in their training dataset.
As summarized in Table I, defenses offering formal bounds achieve low attack accuracies even after many years. Adaptive Tamaraw preserves provable guarantees while introducing clustering and adaptive switching mechanisms to reduce overhead. Crucially, unlike supersequence methods, it remains applicable to websites beyond the training set, making it suitable for real-world, out-of-training scenarios.
III Threat Model
Following prior work [15, 14, 22], we assume a local, passive adversary who can observe all traffic between the client and the guard node in Tor. The adversary cannot alter traffic but can record packet size, timing, and direction to infer the visited webpage. We also assume the adversary is capable of segmenting traffic into individual page loads, and users visit one page at a time. This assumption is standard in WF evaluations to simplify analysis and avoid assumptions that favor the defense.
To carry out the attack, the adversary collects training data by visiting a set of monitored webpages (i.e., those of particular interest) and recording their traffic traces under conditions similar to the target user’s environment. The adversary’s goal is to recognize which of these monitored sites the client is loading.
In this work, we apply two evaluation scenarios for the proposed defense: one where the user’s traffic is limited to webpages seen during the defense’s training phase (in-training webpages), and another where the user can also visit webpages not included in the defense training set (out-of-training webpages). In this latter case, we examine how the defense performs when users browse pages that were not part of the training data used to construct the anonymity sets or train the classifiers. This setting reflects real-world usage, where users may access arbitrary webpages beyond the training set, allowing us to assess the defense’s ability to protect previously unseen traffic patterns.
IV Problem Statement
When a client loads a webpage, it produces a sequence of encrypted packets called a traffic trace:
$f=\bigl{[}(t_{1},d_{1}),(t_{2},d_{2}),\ldots,(t_{N},d_{N})\bigr{]},$
where each packet $f_{i}$ has a timestamp $t_{i}$ and direction $d_{i}\in\{+1,-1\}$ (outgoing/incoming). Since Tor pads packets to a fixed size, the adversary observes only timing and direction.
A defense modifies the trace $f$ to produce a defended trace $f^{\prime}$ by delaying real packets ($t_{i}\mapsto t_{i}+\theta_{i}$) or adding dummies $(t^{\prime},d^{\prime})$, obscuring patterns used by WF attacks.
We evaluate defenses using:
  • Bandwidth Overhead:
    $B(f,f^{\prime})=\frac{|f^{\prime}|-N}{N},$
    the ratio of dummy to real packets.
  • Time Overhead:
    $T(f,f^{\prime})=\frac{t^{\prime}-t_{N}}{t_{N}},$
    the relative delay of the last real packet, where $t^{\prime}$ is the timestamp of the last real packet in the defended trace.
In this paper, we aim to address three limitations of existing WF defenses. First, static padding schedules that ignore real-world traffic diversity are inefficient. Second, supersequence-based defenses struggle in out-of-training settings with unseen websites. Third, defenses that lack provable security guarantees are frequently broken by later attacks. The following expands on each challenge.
1. Excessive Overhead from Static Regularization
Regularization defenses (such as Tamaraw) shape the traffic so that traces appear uniform, minimizing the features an attacker might exploit. However, these methods typically apply a fixed set of parameters statically across all traffic. This static configuration does not account for the inherent variability in traffic patterns across different webpages, which can result in either excessive overhead or insufficient obfuscation in certain cases.
2. Poor Generalization in Supersequence-Based Defenses
Supersequence-based approaches attempt to create anonymity sets by mapping each webpage to a common supertrace. While this strategy can provide strong theoretical guarantees, it is only effective on the set of in-training webpages used during defense construction. If a user’s traffic originates from an out-of-training webpage, there is no reliable mechanism to map the trace to a corresponding supersequence. This restriction greatly limits the applicability of such defenses in real-world deployments. We demonstrate that our method generalizes effectively to out-of-training webpages, making it more suitable for practical scenarios involving previously unseen inputs.
3. Lack of Provable Security Guarantees
Existing defenses have usually been defeated by later-published attacks. In particular, as it can be seen in Table I, state-of-the-art obfuscation defenses (e.g., FRONT, Surakav) often achieve relatively low overhead but have been circumvented by advanced WF attacks. A key limitation of these defenses is the lack of provable security; they do not provide a theoretical upper bound on the adversary’s maximum success rate.
To provide an upper bound on attacker success, we start by analyzing the non-injectivity of a defense. Let $D$ map an input trace $f$ to an output trace $f^{\prime}=D(f)$. For every output, the pre-image is
$D^{-1}(f^{\prime})\;=\;\{f:D(f)=f^{\prime}\}.$
A defense is called $\delta$-non-injective with respect to a given defended trace $f^{\prime}$ if $\bigl{|}D^{-1}(f^{\prime})\bigr{|}\geq\delta$. An attacker who sees $f^{\prime}$ must then guess among at least $\delta$ candidates to infer $f$. However, multiplicity alone is not sufficient for website fingerprinting security: the attacker’s success also depends on how those inputs are distributed across webpages. If one webpage dominates the pre-image, the majority vote still yields a high success rate, even when $\bigl{|}D^{-1}(f^{\prime})\bigr{|}$ itself is large. The relevant quantity is label diversity: the ratio between the total pre-image size and the largest webpage-specific share.
To capture both multiplicity and label diversity, we refine the measure. Let $D^{-1}_{w}(f^{\prime})=\{\,f\in D^{-1}(f^{\prime}):\text{class}(f)=w\}$ be the pre-image subset that originates from webpage $w$. We define the weighted pre-image size
$\tilde{\delta}(f^{\prime})\;=\;\frac{|D^{-1}(f^{\prime})|}{\displaystyle\max_{w}|D^{-1}_{w}(f^{\prime})|}.$ (1)
Intuitively, $\tilde{\delta}(f^{\prime})$ counts how many distinct inputs are merged per majority webpage. If every input in the bucket comes from the same site, the denominator equals the numerator and $\tilde{\delta}(f^{\prime})=1$, so a perfect information attacker will always guess the class of $f^{\prime}$ correctly.
Weighted $\delta$-non-injectivity. We call a defense weighted $\delta$-non-injective with respect to a given defended trace $f^{\prime}$ if $\tilde{\delta}(f^{\prime})\geq\delta$. Since an optimal attacker chooses the majority webpage, its success rate is $\max_{w}|D^{-1}_{w}(f^{\prime})|/|D^{-1}(f^{\prime})|=1/\tilde{\delta}(f^{\prime})$. Thus, a weighted $\delta$-non-injective defense guarantees
$\Pr[\text{attacker succeeds}\mid f^{\prime}]\;\leq\;\frac{1}{\delta}.$
Non-Uniformly Weighted $\delta$-Non-Injectivity
While the above definition bounds the attacker’s success on an individual trace, prior works and subsequent WF studies  [8, 9, 10] focus on the average success rate over the distribution of traces. To capture this, we define the non-uniformly weighted $\delta$-non-injectivity property.
Let $\mathcal{F}^{\prime}$ denote the support of defended outputs, and let $P(f^{\prime})$ be the probability of observing a defended trace $f^{\prime}$ with $f^{\prime}\in\mathcal{F}^{\prime}$. We say the defense is non-uniformly weighted $\delta$-non-injective if this inequality holds:
$\mathbb{E}_{f^{\prime}\in\mathcal{F}^{\prime}}\left[\Pr[\text{attacker succeeds}\mid f^{\prime}]\right]=\mathbb{E}_{f^{\prime}\in\mathcal{F}^{\prime}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right]\leq\frac{1}{\delta}.$
This definition implies that the attacker’s success rate, on average, remains below the inverse of $\delta$. This choice aligns with previous work on website fingerprinting defenses, which report attacker performance in terms of average success rate over the trace distribution, i.e., they evaluate non-uniform security. Uniform security, while stricter, is rarely achieved in practice and not the focus of most deployed defenses. We therefore adopt the non-uniform formulation to remain consistent with established evaluation methodology [10].
V Design Details
V-A Overview
In this work, we introduce the first website fingerprinting defense that performs adaptive parameter selection while providing a formal upper bound on attacker success. Our design is motivated by three goals: (i) replacing static regularization configurations with traffic adaptation, (ii) generalizing to webpages not seen during training, and (iii) ensuring provable security guarantees.
We do so using a global-to-local defense strategy. Our defense assumes no knowledge of the page under load, so it always starts with a global strategy based on a regularization defense (such as BuFLO [21], Tamaraw [10], or RegulaTor [12]). We begin each page load using a set of global parameters that apply conservative padding and scheduling to protect early traffic without knowing the destination. After enough of the trace has been observed to confidently identify an anonymity set, we switch to a more efficient set of padding parameters tailored to that set. This two stage strategy ensures strong privacy guarantees during the early ambiguous phase while reducing overhead in the remainder of the trace.
To concretely instantiate this global-to-local strategy, we build on the Tamaraw defense, leading to what we call Adaptive Tamaraw.
V-B Adaptive Tamaraw
To evaluate the general framework introduced in this paper, we use a concrete, well-studied regularization defense in our global-to-local strategy: Tamaraw [10]. Among existing regularization-based defenses, Tamaraw stands out as the only one that offers a formal, information-theoretic bound on an adversary’s success rate. As shown in Table I, it also yields the lowest attacker accuracy compared to other defenses.
Classic Tamaraw transmits fixed-size cells at two constant rates: $\rho_{\text{out}}$ seconds between successive upstream cells and $\rho_{\text{in}}$ seconds between successive downstream cells. Padding continues in each direction until its cell count reaches the next multiple of a hyperparameter $L$. Since $(\rho_{\text{out}},\rho_{\text{in}})$ are fixed and identical across all traces, the resulting shapes of the defended traces are uniform in timing and rate. The only distinguishing feature that may remain between traces is their total length: specifically, which multiple of $L$ they are padded to.
Tamaraw requires a single pair of padding parameters $(\rho_{\text{out}},\rho_{\text{in}})$ to be applied uniformly across all webpages. This global configuration leads to inefficiencies, as some webpages end up causing far more dummy traffic than necessary. As noted by [10], the standard method for configuring Tamaraw involves fixing the $L$ and then performing a grid search over candidate $(\rho_{\text{out}},\rho_{\text{in}})$ pairs. Each pair yields a specific tuple of bandwidth and time overhead, and the set of resulting points is filtered to retain only the Pareto-optimal ones (those that are not dominated in both overhead dimensions).
In Adaptive Tamaraw, we start each trace with these global parameters, and then switch to local padding parameters when possible. This strategy is governed by three processes that will be described in the subsequent subsections.
  1. 1.
    We analyze the webpages in the training set of the defense to identify representative traffic patterns, which we call Intra-Webpage Pattern Detection (Section V-C).
  2. 2.
    These traffic patterns are grouped together in Anonymity Set Generation (Section V-D), wherein local parameters are identified for each anonymity set.
  3. 3.
    Early Anonymity Set Detection (Section V-E) tells the defense when to switch from global to local parameters based on packets of the observed live trace.
V-C Intra-Webpage Pattern Detection
Webpages do not generate a single, uniform traffic pattern; rather, they produce a number of patterns. This can be due to: dynamically generated content such as personalized recommendations [23], advertisements that can differ significantly in size and frequency [24], and variations in localization tailored for different countries [25]. As an illustrative example, Figure 1 shows four traces from the same webpage in the Sirinam et al. dataset [6], visualized using the Traffic Aggregation Matrix (TAM) [8]. The two rows highlight two distinct recurring structures, highlighting that even a single page can yield multiple characteristic patterns.
(a) Pattern 1
(b) Pattern 2
Figure 1: Visualization of the TAMs of four traces from the same page, in the dataset obtained by [6]. Each TAM divides the first 1000 time-slots (80 ms per slot) into rows for outgoing packets (blue, positive values) and incoming packets (red, negative values); the height of each bar records the packet count in that slot. The first two traces (top row) exhibit a very similar pattern, while the next two traces (bottom row) share a distinct structure. This highlights the possibility of multiple recurring traffic patterns within a single page.
As a result of these observations, aggregating all traces from a single webpage into one homogeneous profile is inefficient and often leads to over-padding and degraded performance. However, many existing defenses construct anonymity sets at the webpage level, assuming that all traces from the same page behave similarly [2, 16, 22]. Instead, our approach aggregates traffic at the pattern level to form smaller, more homogeneous groups.
To obtain distinct traffic patterns within each webpage, we first represent each network trace as a time series using the TAM representation introduced in  [8]. In our TAM representation, we divide the total page load time into a fixed number of time slots and record the number of incoming and outgoing packets in each slot. This results in a two-dimensional time series, where each trace can be viewed as a bivariate sequence capturing the dynamics of incoming and outgoing traffic over time.
With traces represented as TAM time series, we cluster the traces of each webpage individually to isolate recurring traffic patterns. We refer to the resulting clusters as intra-webpage clusters (i.e., for the remainder of this section, when we refer to intra-webpage clusters, we mean clusters obtained by grouping traces originating from a single webpage). To perform this task, we adopt the Cluster Affinity Search Technique (CAST) [26]. CAST does not require us to specify the number of clusters in advance; instead, it uses an affinity threshold $T$ (which is a hyperparameter) to determine membership. The affinity of a time series $F_{x}$ with respect to a particular intra-webpage cluster $\mathcal{S}$ is defined as follows:
$a_{\mathcal{S}}(F_{x})\;=\;\frac{1}{\lvert\mathcal{S}\rvert}\sum_{F_{y}\,\in\,\mathcal{S}}A(F_{x},F_{y}).$
where $A(F_{x},F_{y})$ is the similarity between time series $F_{x}$ and $F_{y}$, and $\lvert\mathcal{S}\rvert$ is the size of the intra-webpage cluster. CAST evolves a cluster by repeatedly selecting the unassigned time series with the highest affinity to the cluster:
  • If its affinity exceeds the fixed threshold $T$, the series is added to the intra-webpage cluster.
  • Otherwise, the algorithm checks if any existing member in the intra-webpage cluster has an affinity falling below the threshold; if so, that low-affinity member is removed.
  • If no member meets the removal criterion, the intra-webpage cluster is deemed complete (closed), and CAST proceeds to form a new intra-webpage cluster with the remaining data.
Our preliminary evaluations revealed that for traces in a single webpage, CAST generated several large intra-webpage clusters, along with a long tail of smaller ones. Small intra-webpage clusters are not suitable, as our goal is to regularize all traces within an intra-webpage cluster to appear identical; if an intra-webpage cluster contains too few traces, the defense incurs the overhead of padding and shaping with minimal privacy benefit. To be more precise, the attacker’s uncertainty remains low despite the added cost, resulting in an inefficient overhead-to-privacy tradeoff. To produce better clusters, we introduced four modifications to CAST, which we will describe below.
Affinity Computation. As we mentioned, the affinity of a time series $F_{x}$ with respect to a candidate cluster $\mathcal{S}$ is computed as the average of the pairwise similarity scores between $F_{x}$ and each member $F_{y}\in\mathcal{S}$. Following prior work [27, 28], the similarity score $A(F_{x},F_{y})$ is computed using an exponential function of the squared Euclidean distance:
$A(F_{x},F_{y})\;=\;\exp\!\Bigl{(}-\frac{d^{2}(F_{x},F_{y})}{\sigma^{2}}\Bigr{)},$
where $\sigma$ is a bandwidth hyperparameter that governs the sensitivity of the similarity function. Previous studies [29] have shown that a global $\sigma$ can fail to capture variations in data density.
To address this, we adopt a local scaling approach inspired by self tuning spectral clustering [30, 31], in which each time series $F_{x}$ is assigned its own scale parameter $\sigma_{x}$, typically set to the distance between $F_{x}$ and its $K$-th nearest neighbor. The similarity is then computed as:
$A(F_{x},F_{y})\;=\;\exp\!\Bigl{(}-\frac{d^{2}(F_{x},F_{y})}{\sigma_{x}\,\sigma_{y}}\Bigr{)}.$
Cleaning Step. After forming the initial clusters, we perform a final pass. For each point $c$ in each cluster $\mathcal{C}_{i}$, we compute its affinity to every other cluster $\mathcal{C}_{j}$. If we find a cluster $\mathcal{C}_{j}$ in which the point $c$ achieves a strictly higher affinity than in its current cluster $\mathcal{C}_{i}$, we remove $c$ from $\mathcal{C}_{i}$ and reassign it to $\mathcal{C}_{j}$. This process is repeated until no point changes its cluster membership.
Post-Processing Step. To limit the number of clusters produced by our modified CAST algorithm, we adopt a post-processing procedure inspired by [32]. Each cluster consists of a set of TAM-based trace representations. We measure its cut and volume using the similarity score $A$ introduced earlier:
$\mathrm{cut}(\mathcal{C})\;=\;\sum_{f\in\mathcal{C}}\sum_{f^{\prime}\notin\mathcal{C}}A(f,f^{\prime}),\quad\mathrm{vol}(\mathcal{C})\;=\;\sum_{f\,,\,f^{\prime}\in\mathcal{C}}A(f,f^{\prime}).$
The expansion ratio $\phi(\mathcal{C})=\mathrm{cut}(\mathcal{C})/\mathrm{vol}(\mathcal{C})$ is low when traces inside $\mathcal{C}$ are highly similar while remaining well separated from traces outside.
We then apply an iterative post-processing routine: as long as the number of clusters exceeds a predefined threshold, we sort the clusters by size, select the smallest cluster, and merge it with a neighboring cluster such that the resulting partition minimizes the largest expansion ratio among all clusters [33].
Dynamic Affinity Threshold. Finally, rather than using a fixed affinity threshold for all traces, we compute it in a data-driven manner. We first calculate the global mean similarity across all pairs of traces:
$T\;=\;\frac{1}{n^{2}}\sum_{x<y\leq n}A(F_{x},F_{y}),$
where $A(F_{x},F_{y})$ is the pairwise similarity between traces $x$ and $y$, and $n$ is the total number of traces. This global average $T$ serves as a baseline for typical similarity values within the dataset. We then set the CAST affinity threshold to $T$ (i.e., the minimum similarity required to add a new trace to an intra-webpage cluster). This approach ensures that the threshold automatically adapts to the overall range of similarities in the dataset, rather than being a manually tuned constant hyperparameter, thereby improving the robustness of the resulting clusters.
We apply the modified CAST algorithm separately to each webpage in the dataset to extract a set of characteristic traffic patterns. These extracted patterns form the foundation for all subsequent phases. The complete pseudocode for this phase is given in Appendix D.
V-D Anonymity Set Generation
Once we have extracted distinct traffic patterns from each webpage, as described in Section V-C, we proceed to cluster these patterns into anonymity sets. This phase organizes structurally similar patterns, regardless of their originating web pages, into groups that will later be reshaped using a shared regularization policy. In contrast to prior work that clusters at the webpage level [2, 16, 22], this fine-grained grouping reduces overhead and improves generalization.
Our anonymity sets are constructed to satisfy two key properties that underlie the provable security guarantee offered by our defense:
  • k-anonymity: Each anonymity set must contain at least $k$ distinct traffic patterns [34]. In this way, even if an adversary correctly identifies the anonymity set to which a trace belongs, they are left with at least $k$ possibilities, bounding the probability of correctly inferring the true pattern to at most $1/k$.
  • l-diversity: The patterns within each anonymity set must originate from at least $l$ different webpages [19]. This criterion ensures that even if the anonymity set is exposed, the attacker cannot infer the correct webpage simply because all patterns come from a single source.
For this phase, we adopt the $k$-anonymity-based clustering algorithm introduced in Palette [22] as our base algorithm. In Palette, anonymity sets are generated at the webpage level. The algorithm first computes the Traffic Aggregation Matrix for each trace of a webpage and then aggregates these into a super-matrix by an element-wise maximum. These super-matrices are then used as the elements (i.e., the representatives of each webpage) for a $k$ anonymity-based clustering algorithm [34], with the similarity between two webpages measured by computing the Euclidean distance between their corresponding super-matrices. The clustering itself follows a greedy strategy: it constructs one anonymity set at a time by first selecting an initial webpage and then repeatedly adding the webpage whose super-matrix is closest, based on Euclidean distance, to the current super-matrix of the partially constructed set. Once $k$ webpages have been assigned, the process repeats to form the next set. The output is a list of anonymity sets, each containing at least $k$ webpages. We adopt the same $k$ anonymity-based clustering algorithm as used in Palette, but with two key modifications:
Pattern-level clustering. Our method operates at the pattern level rather than the webpage level. As explained in Section V-C, a single webpage can produce multiple distinct traffic patterns due to various reasons; clustering at this finer granularity yields more homogeneous groups. As an illustrative example, Figure 2 compares the impact of clustering at the webpage level versus the pattern level. For each fixed value of $k$, we conduct two separate experiments. In the first, we compute one supermatrix per webpage by taking the element-wise maximum over all TAMs from that webpage, and then apply the clustering algorithm from Palette to these webpage-level supermatrices. In the second, we repeat the process at the pattern level by computing one supermatrix per extracted traffic pattern and clustering those instead, again using Palette. In both cases, after clustering, we treat each cluster’s supermatrix as the defended version of all traces in the cluster and compute the bandwidth overhead across all the traces. This provides a proxy for how well the clustering captures homogeneity. As shown in the figure, clustering at the pattern level consistently yields lower overhead.
Figure 2: Comparison of website-level vs. pattern-level clustering. For each value of $k$, clustering is applied to supermatrices constructed at the website or pattern level, and the resulting average bandwidth overhead is measured. Pattern-level clustering consistently results in lower overhead, especially as $k$ increases, indicating that clustering finer-grained traffic patterns captures homogeneity more effectively than aggregating at the website level.
Diversity-aware distance metric. Instead of computing the Euclidean distance between super-matrices, we define a distance function $d$ that directly captures attacker success after regularization. Given an anonymity set $C$ currently under construction and a candidate pattern $p$ (i.e., an intra-webpage cluster), we evaluate the impact of merging them by computing the average-case attacker success rate over the combined set $C^{\prime}=C\cup p$.
For each pair of regularization parameters $(p_{\text{in}},p_{\text{out}})$ in a predefined grid $\mathcal{P}$, we apply Tamaraw to all traces in $C^{\prime}$ and compute the non-uniform attacker accuracy:
$\bar{A}\bigl{(}C^{\prime};p_{\text{in}},p_{\text{out}}\bigr{)}\;=\;\sum_{\ell}\frac{|C^{\prime}_{\ell}|}{|C^{\prime}|}\cdot\frac{\displaystyle\max_{w}\bigl{|}\{\,t\in C^{\prime}_{\ell}:\mathrm{site}(t)=w\}\bigr{|}}{|C^{\prime}_{\ell}|},$
where $C^{\prime}_{\ell}$ denotes the subset of traces in $C^{\prime}$ whose regularized lengths equal $\ell$. This measures the expected success rate of an optimal attacker who predicts the most frequent webpage label in each length bucket.
The distance $d(C,p)$ is then defined as the average attacker accuracy over the entire grid:
$d(C,p)\;=\;\frac{1}{|\mathcal{P}|}\sum_{(p_{\text{in}},p_{\text{out}})\in\mathcal{P}}\bar{A}\bigl{(}C^{\prime};p_{\text{in}},p_{\text{out}}\bigr{)}.$
This formulation implicitly enhances $l$-diversity, since adding a pattern from a new webpage typically reduces the attacker’s success rate based on the webpage in the majority, thus lowering $d(C,p)$. We provide empirical validations for this attribute in Appendix B-C. During clustering, we therefore always merge the candidate $p$ that minimizes $d(C,p)$, encouraging both homogeneity in regularized shape and diversity in origin. Algorithm 1 provides detailed pseudocode of our proposed anonymity set generation algorithm.
Algorithm 1 Pattern-Level Anonymity Set Generation
Input:
$P=\{p_{1},\dots,p_{N}\}$: all extracted patterns
$k$: anonymity parameter
$d(S,p)$: distance function between anonymity set $S$ and pattern $p$
Output:
$\mathcal{S}=\{S_{1},\dots,S_{m}\}$: final anonymity sets
[…]
3:Step 1: Seed the first set
4:Choose any $p_{\text{seed}}\in U$
[…]
6:Step 2: Build additional sets
7:for $i=1$ to $\left\lfloor\frac{|P|}{k}\right\rfloor$ do
8:  if $|S_{i}|<k$ then
9:   while $|S_{i}|<k$ and $U\neq\varnothing$ do
10:     $p^{\star}\leftarrow\arg\min_{p\in U}d(S_{i},p)$
11:     $S_{i}\leftarrow S_{i}\cup\{p^{\star}\}$; $U\leftarrow U\setminus\{p^{\star}\}$
12:   end while
13:  end if
14:  if $U\neq\varnothing$ then
15:   $p_{\text{new}}\leftarrow\arg\max_{p\in U}\sum_{S\in\mathcal{S}}d(S,p)$
[…]
17:  end if
18:end for
19:Step 3: Assign remaining patterns
20:for all $p\in U$ do
21:  $j\leftarrow\arg\min_{S\in\mathcal{S}}d(S,p)$
22:  $S_{j}\leftarrow S_{j}\cup\{p\}$
23:end for
24:return $\mathcal{S}$
The two previous phases (pattern extraction and anonymity set generation) comprise the offline training stage of our framework. A high-level illustration of this process is shown in Figure 3, which depicts how traffic traces from different webpages are transformed into anonymity sets based on structured patterns.
Figure 3: High–level workflow of the first two phases of our defense. 1. Pattern extraction. For each webpage in the training dataset, we group its traces into a small number of stable, recurring traffic patterns (dashed boxes), reflecting variability due to CDNs, localization, and user behavior. 2. Anonymity set construction. The extracted patterns are then clustered across different webpages to form anonymity sets. In this example, each set contains at least $k=3$ distinct patterns originating from at least $l=2$ different webpages, thereby satisfying $k$-anonymity and $l$-diversity. A lightweight, cluster-specific regularization schedule is precomputed for each set.
V-E Early Anonymity Set Detection
As we mentioned, our defense follows a global-to-local regularization strategy. Every new page load is protected by singular global regularization parameters until enough packets have arrived to identify the anonymity set that best matches the live prefix; once the set is known, we switch to lighter, local regularization parameters.
For the global-to-local switch, deciding which set to choose and when it is safe to switch is an instance of early time-series classification, a task where the goal is to make accurate predictions based on incomplete sequences, as early as possible. Early classification has been studied extensively in latency-critical domains such as medical diagnosis [35, 36] and industrial process monitoring [37, 38]. In this work, we adopt the ECDIRE framework [39], which has been designed for reliable early classification of time series. ECDIRE learns, for every class (in our case anonymity set), the earliest safe timestamp at which the class can be predicted with at least a user-selected fraction $\alpha$ of the accuracy that would be achieved on complete traces. During deployment the classifier is queried only at those safe timestamps, which removes unnecessary tests on very short prefixes and guarantees that a decision is issued as soon as the chosen confidence threshold is met.
For every anonymity set $\mathcal{S}$ we compute a safe timestamp $\tau_{\mathcal{S}}$ as the earliest prefix length (i.e., time) $t$ at which a validator (described below) can distinguish $\mathcal{S}$ from all other sets with accuracy $\geq\alpha\,A_{\mathcal{S}}^{\text{full}}$, where $A_{\mathcal{S}}^{\text{full}}$ is the accuracy of the validator on full traces of $\mathcal{S}$. We never test $\mathcal{S}$ before $\tau_{\mathcal{S}}$; conversely, if the trace is not accepted at $\tau_{\mathcal{S}}$, $\mathcal{S}$ is never considered again for that trace. This “single‑shot” rule forces every trace that joins $\mathcal{S}$ to switch defenses at the same fixed time, closing a potential timing side channel.
In the original ECDIRE algorithm [39], the training process initially involves training a separate classifier at each timestep. Once training is complete, a post-processing step identifies the safe prediction timestamps, the points in time where early classification is both accurate and confident. The final model then retains only the classifiers corresponding to these selected timestamps, discarding the rest. In our case, WF classifiers are deep networks, so to train and retain a full model for each timestamp would be prohibitively expensive in time and memory. Instead we factor the prediction task into a single deep backbone plus a collection of lightweight per-site models as follows.
  1. 1.
    Stage A – webpage predictor (Holmes). We train one Holmes network [40], a spatial–temporal CNN encoder learned with supervised contrastive loss so that partial traces embed close to their full counterparts. At inference time Holmes outputs the most likely webpage $w$ given an incomplete input trace.
  2. 2.
    Stage B – Pattern predictor (per‑site k‑fingerprinting). For each webpage $w$ and each time stamp $t$ of interest, we train a small k fingerprinting (kFP) random forest classifier [4], denoted $\text{kFP}_{w}^{(t)}$, using the traces of the site truncated at time $t$ and the labels of each trace being the number of the cluster intra-site (i.e. pattern) of that trace. kFP requires only a modest number of examples, making it well-suited for scenarios where per-site data is limited.
Given an incoming trace prefix of length $t$, the prediction process proceeds as follows:
  1. 1.
    Holmes predicts the most likely webpage $w$.
  2. 2.
    The corresponding kFP model $\text{kFP}_{w}^{(t)}$ identifies the pattern $p$.
  3. 3.
    By construction, the pair $(w,p)$ uniquely determines an anonymity set $\mathcal{S}_{w,p}$. If $t=\tau_{\mathcal{S}_{w,p}}$ (i.e., the current timestamp is the safe prediction timestamp for that anonymity set), we immediately switch to that set’s lighter parameters; otherwise, we continue applying the global regularization schedule.
After training the single Holmes model and all $\text{kFP}_{w}^{(t)}$ classifiers across timestamps, we determine a safe timestamp $\tau_{\mathcal{S}}$ for each anonymity set $\mathcal{S}$, indicating the earliest point at which a confident classification can be made. For each webpage $w$, we keep only the $\text{kFP}_{w}^{(t)}$ models corresponding to its safe timestamps. The Holmes model is site-agnostic and trained once, then reused across all traces and timesteps. This design preserves ECDIRE’s early-decision capability while reducing computational cost by avoiding deep model training at every time step. Figure 4 illustrates this switching process and its alignment with safe timestamps.
Figure 4: Illustration of early anonymity set prediction and parameter switching. An incoming trace is initially regularized using the global parameters $(p_{\text{in}}^{g}=0.012,p_{\text{out}}^{g}=0.04)$ in this example. At each predefined safe timestamp, the classifier attempts to assign the trace to one of the candidate anonymity sets. In the first attempt, the classifier predicts $S_{1}$, but no switch occurs because $S_{1}$ is not valid at that timestamp. At the second safe timestamp, the trace is matched with $S_{2}$, which is an acceptable set at that point. This triggers a transition to $S_{2}$’s per-set lighter parameters $(p_{\text{in}}^{S_{2}}=0.06,p_{\text{out}}^{S_{2}}=0.08)$, which are applied for the rest of the connection. Each anonymity set is tied to a unique safe timestamp, so switching can occur at most once.
V-F Security Bound for Adaptive Tamaraw
Despite transitioning from a fixed global defense to lighter, AS-specific configurations after a safe prediction point, Adaptive Tamaraw retains a provable security guarantee. Specifically, we show that the defense remains non-uniformly weighted $\delta$-non-injective, ensuring that the attacker’s average success probability is formally bounded.
Theorem V.1 (Global Non-Uniformly Weighted $\delta$–Non-Injectivity).
Let $\mathcal{S}$ be the set of anonymity sets constructed in Section V-D. For a fixed global regularization parameter pair $(p_{\text{in}},p_{\text{out}})$, let $\bar{A}(S_{i};p_{\text{in}},p_{\text{out}})$ denote the expected attacker success rate over anonymity set $S_{i}\in\mathcal{S}$, as defined in Section VI-D (Eq. 5). Define the global injectivity parameter $\delta$ as:
$\frac{1}{\delta}=\mathbb{E}_{S_{i}\sim\mathcal{S}}\left[\bar{A}(S_{i};p_{\text{in}},p_{\text{out}})\right].$
Then Adaptive Tamaraw is non-uniformly weighted $\delta$-non-injective, and the attacker’s average success probability is bounded by:
$\Pr[\text{success}]\;\leq\;\frac{1}{\delta}.$
The formal proof of this result is presented in Appendix C.
This result is intuitive: the attacker’s overall bounded success rate is the average across the bound for each combination of anonymity set and trace length. Any two traces from the same set and of the same length are made indistinguishable by the defense. Consequently, the attacker’s optimal strategy in an anonymity set reduces to guessing the most common webpage in each such length-matched group, which limits their average success rate captured by $\bar{A}(S_{i})$.
VI Experimental Evaluation
In this section, we empirically evaluate the effectiveness, efficiency, and generalizability of Adaptive Tamaraw. Our goal is to demonstrate that the proposed defense reduces bandwidth and latency overhead while maintaining provable bounds on attacker success rates, even under strong adversarial models. Section VI-A describes the experimental setup, including datasets, hyperparameters, and model architectures. Section VI-B evaluates performance on in-training webpages, where all defended sites were included in the training phase. Section VI-C extends the analysis to out-of-training webpages, assessing the defense’s robustness against previously unseen traffic patterns. Finally, Section VI-D verifies the tightness of the theoretical bounds established in Section V-F, comparing them against observed attack accuracies.
VI-A Experimental Setup
Dataset
We conduct our experiments using the widely adopted Tor-based website fingerprinting dataset originally collected by [6], which has become a standard benchmark in recent WF research, including works such as [8, 22, 41, 42]. The dataset comprises two parts: a closed-world set and an open-world set. The closed-world subset includes 1,000 traffic traces for each of 95 monitored webpages, totaling 95,000 labeled instances. All traces were collected via real Tor connections and reflect authentic network conditions including congestion, jitter, and Tor circuit variability.
Parameter Grid for Adaptive Tamaraw
Tamaraw is parameterized by three key hyperparameters: the upstream and downstream padding intervals $(\rho_{\text{out}},\rho_{\text{in}})$, and the length bucket parameter $L$: all traces are padded to multiples of $L$. A higher $L$ increases the likelihood that multiple defended traces will collapse to the same length, reducing fingerprintability. We consider three values for $L$: $\{100,500,1000\}$.
For each $L$, we explore a grid of candidate values for $\rho_{\text{in}}$ and $\rho_{\text{out}}$. Given that incoming traffic typically arrives more frequently than outgoing traffic during webpage loading [10, 12], we sweep $\rho_{\text{in}}$ in the range $[0.001,0.006]$ and $\rho_{\text{out}}$ in the range $[0.005,0.21]$. We retain only the Pareto-optimal configurations (those for which no other configuration achieves strictly lower bandwidth and time overhead simultaneously [10]). This selection yields 33 unique $(\rho_{\text{in}},\rho_{\text{out}})$ pairs that span the optimal trade-off frontier between efficiency and security. A detailed report of these configurations is provided in Appendix B-A.
Model Setup
As discussed in Section V-E, our early anonymity set detection relies on a two-stage pipeline: a Holmes model for webpage-level prediction, followed by a lightweight k-fingerprinting (kFP) model for identifying fine-grained traffic patterns. Full details on model architectures are provided in Appendix A.
Hyperparameter Configuration
To ensure optimal performance and fair evaluation, we conducted a grid search for each hyperparameter over its respective search space. Table II summarizes the key hyperparameters used in Adaptive Tamaraw and its underlying components. The first three parameters ($k$, Max Intra-Webpage Patterns, and $\alpha$) were selected to minimize overhead while preserving security guarantees. The remaining model-specific hyperparameters for the Holmes and kFP classifiers were chosen to maximize prediction accuracy in the early anonymity set detection phase.
TABLE II: Hyperparameters for Adaptive Tamaraw with grid search spaces.
Parameter Value Search Space
$K$ (local scaling) 7 1–10
Max Intra Webpage Patterns 6 2–8
$\alpha$ (ECDIRE) 0.9 {0.6,0.7,0.8,0.9, 1}
Holmes lr 5e-4 {5e-6,5e-5,5e-4,5e-3}
TAM Time Slot 80ms {40ms, 80ms, 120ms }
Holmes batch 256 {32,64,128,256}
Holmes epochs 80 fixed
VI-B In-Training Evaluation
In the in-training setting, we evaluate the effectiveness of our defense in protecting webpages that were explicitly included in the defense’s training dataset. Specifically, we focus on the 95 monitored webpages from the closed-world portion of the dataset introduced in Section VI-A. These sites serve as our set of protected targets. Following prior work [22, 8], we partition the dataset into training, validation, and testing sets using an 8:1:1 ratio. At a high level, our full defense pipeline proceeds as follows. We begin by applying the modified CAST algorithm V-C to the training traces of each webpage to extract recurring traffic patterns. These patterns are then clustered using the $k$-anonymity-based algorithm introduced in Section V-D to obtain our anonimity-sets (we perform the experiments by varying the minimum required anonymity set size $k$ from 2 to 30). For each of the obtained anonymity-sets, we identify a safe timestamp using early time-series classification as described in Section V-E, allowing us to switch from a global Tamaraw configuration to a cluster-specific configuration as early as possible. Finally, we compare the performance against applying the classical Tamaraw to the traces.
For a fixed global configuration $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$ selected from the Pareto-optimal grid introduced in Section VI-A, we determine cluster-specific regularization parameters for each anonymity set. Specifically, we sweep over the full grid of $(\rho_{\text{in}},\rho_{\text{out}})$ pairs and select the configuration that achieves lower bandwidth and time overheads on the traces within the anonymity set, compared to $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$. This ensures that the local configuration provides a clear improvement over the global baseline for the assigned subset of traces.
We also train the Holmes model on the same training traces to perform early webpage prediction, and independently train a set of kFP classifiers for each webpage to support fine-grained pattern prediction. Using the ECDIRE procedure (described in Section V-E), we derive a safe prediction time for each $(\text{webpage},\text{pattern})$ pair. We set $\alpha=0.9$, meaning a prediction is considered safe once it reaches 90% of the classification accuracy achieved when the classifier is trained on complete traces. This value was selected empirically: we evaluated a range of $\alpha$ values from 0.5 to 1, and compared the resulting overhead improvements. The experiments showed that $\alpha=0.9$ yielded the best overhead improvements compared to the original tamaraw.
We proceed to evaluate the performance of Adaptive Tamaraw using the held-out test set. For each test trace $i$, and for each global Pareto regularization pair $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$, we first apply Tamaraw with the global parameters and record the resulting bandwidth and time overheads. Next, we evaluate the same trace under the complete adaptive Tamaraw strategy: we begin with $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$ and switch to the specific configuration of the anonymity set $(\rho_{\text{in}}^{\mathcal{S}_{i}},\rho_{\text{out}}^{\mathcal{S}_{i}})$ once the predicted anonymity set becomes identifiable according to its safe prediction time. We then record the resulting overheads. We repeat this procedure for all the 33 global Pareto regularization pairs $(\rho_{\text{in}}^{g},\rho_{\text{out}}^{g})$.
Table III presents average total overhead (combined time and bandwidth, averaged over all global regularization pairs) for static Tamaraw versus Adaptive Tamaraw at various cluster sizes $k$, across three bucket lengths $L\in\{100,500,1000\}$. Adaptive Tamaraw achieves notable reductions across all settings, particularly for small $k$. For example, when L=1000L{=}1000, Adaptive Tamaraw reduces the average overhead from 459% (Tamaraw) to 360% at k=2k{=}2, a drop of 99%. Even at k=30k{=}30, the reduction remains significant at 60%. For smaller $L$, the relative savings are smaller but still meaningful: at L=100L{=}100 and k=2k{=}2, the overhead drops from 243% to 205% (38% decrease), while at L=500L{=}500, the drop is from 299% to 258% (41%).
The choice of $k$ controls the defense’s effectiveness: smaller values result in a looser bound on attacker accuracy (fully explored in Section VI-D) while producing the smallest overhead. Interestingly, Adaptive Tamaraw results in higher overhead savings with smaller $k$, i.e. when the anonimity sets cover less intra-webpage patterns. An intuitive explanation is that in high $k$ settings, each anonymity set is large and therefore does not benefit as much from localized, set-specific adaptive parameters. More comprehensive results across a broader range of $k$ values are provided in Appendix B-B.
TABLE III: Average total overhead (time + bandwidth) comparison between static Tamaraw and Adaptive Tamaraw across multiple cluster sizes $k$ and bucket lengths $L$. Each entry shows the absolute overhead percentage, with the difference from Tamaraw in parentheses. Adaptive Tamaraw consistently reduces overhead, especially for small values of $k$, while maintaining security guarantees.
L Tamaraw (Baseline) Adaptive Tamaraw
k = 2 k = 7 k = 30
100 243.0% 205.0% (-38.0) 217.0% (-26.0) 235.0% (-8.0)
500 299.0% 258.0% (-41.0) 270.0% (-29.0) 283.0% (-16.0)
1000 459.0% 360.0% (-99.0) 379.0% (-80.0) 399.0% (-60.0)
Per-Trace Analysis. Beyond aggregate averages, we examine how Adaptive Tamaraw impacts overhead on a per-trace basis. To this end, we selected a representative global Pareto configuration, $(\rho_{\text{in}}=0.012,\rho_{\text{out}}=0.04)$, which was also reported in the original Tamaraw paper [10], and offers a balanced trade-off between bandwidth and time overhead. We applied both static Tamaraw and Adaptive Tamaraw (with $k=7$ and $L=100$) to each test trace using this configuration as the global parameters. The resulting average bandwidth and time overheads for static Tamaraw were 106% and 53%, respectively. With Adaptive Tamaraw, these dropped to 95% and 47%, corresponding to an overall overhead reduction of approximately 17%. We measured the difference in total overhead between the two methods for each trace and plotted the distribution.
Figure 5 shows the resulting distribution of per-trace total overhead savings. Positive values indicate cases where Adaptive Tamaraw yielded lower overhead than static Tamaraw, while negative values represent regressions. Although some traces exhibit higher overhead under adaptation, the overall trend is clearly beneficial: the average saving is positive and approximately 10% of traces experience savings exceeding 100%, even reaching up to 500%. These results demonstrate Adaptive Tamaraw’s ability to reduce overhead, particularly significant for traces that would otherwise be heavily penalized by a one-size-fits-all padding strategy.
Furthermore, to assess how reliably Adaptive Tamaraw identifies the correct anonymity set at runtime, we evaluated the performance of the combined Holmes + kFP classifier used in the ECDIRE procedure. On average, the correct anonymity set was identified in 81% of test traces. In 10% of cases, no set was chosen, resulting in the defense completing using the original Tamaraw parameters. The remaining 9% of traces were assigned to an incorrect anonymity set.
This per-trace analysis underscores a key advantage of adaptation: by tailoring the padding schedule to individual traffic characteristics as soon as sufficient information becomes available, our approach delivers substantial efficiency gains in a fine-grained and targeted manner.
Figure 5: Distribution of per-trace overhead savings achieved by Adaptive Tamaraw over static Tamaraw for one representative global padding configuration: $(\rho_{\text{in}}=0.012,\rho_{\text{out}}=0.04)$, with $k=7$ and $L=100$.The red vertical line at 0% indicates the point where both methods incur equal overhead; values to the right indicate savings from adaptation, and those to the left indicate higher cost.
VI-C Out-of-Training Evaluation
In the out-of-training setting, our goal is to assess how well Adaptive Tamaraw generalizes to website traffic that was not included during the training phase. This scenario is particularly important, as prior anonymity set–based defenses [16, 2, 22] are typically constrained to protect only in-training webpages, limiting their applicability in real-world browsing environments where users may visit previously unseen pages.
We split the closed-world dataset, which comprises 95 webpages, into two disjoint subsets of webpages. One subset (containing approximately half of the webpages) serves as the training set for the defense, while the other subset is reserved for the out-of-training evaluation, simulating previously unseen webpages.
For this experiment, we instantiate Adaptive Tamaraw with $k=7$ and $L=100$, a configuration that strikes a practical balance between efficiency and protection. The choice of $L=100$, inspired by prior work [10], provides reasonable security guarantees while avoiding the excessive overhead associated with larger values. Meanwhile, $k=7$ offers an effective trade-off: it yields notable reductions in bandwidth and time overhead (Table III) and maintains low prediction accuracy within anonymity sets (Section VI-D) (i.e., under this setting, the maximum theoretical success rate of any attacker remains below 50%).
We follow the same training procedure as in the in-training setting, including pattern extraction, anonymity set generation, classifier training, and safe time on the first half of webpages. During evaluation of the testing traces, the defense initially applies the global Tamaraw parameters from Section VI-A and then attempts to assign the trace to one of the anonymity sets, constructed exclusively from the training webpages, at any of the precomputed safe times. If such an assignment is possible, the defense switches to the corresponding anonymity set–specific parameters for the remainder of the trace.
To compare the efficiency of different defense strategies under practical constraints, we fix a series of time overhead thresholds and, for each threshold, identify the configuration that yields the lowest bandwidth overhead while keeping the time overhead below the given limit. We report this value for three scenarios: (i) the original Tamaraw defense, (ii) Adaptive Tamaraw on out-of-training webpages, and (iii) Adaptive Tamaraw on in-training webpages. We threshold time overhead as it represents the maximum tolerance a user can be expected to accept for a defense: time delays directly affect user experience while bandwidth affects network load.
TABLE IV: Comparison of bandwidth overhead for Tamaraw and Adaptive Tamaraw under various fixed time overhead constraints. Results are reported for both in-training and out-of-training settings of Adaptive Tamaraw. Values in parentheses represent percentage reductions relative to the original Tamaraw. Adaptive Tamaraw consistently lowers both minimum and average bandwidth overhead, with the in-training variant achieving the largest reductions.
Fixed Time Overhead Minimum Bandwidth Overhead (%) Average Bandwidth
<< 10% << 45% << 125% << 250% Overhead (AUC)
Tamaraw 279 124 83 66 82
Adaptive Tamaraw (out-of-training) 277 (-2) 119 (-5) 80 (-3) 64 (-2) 79 (-3 )
Adaptive Tamaraw (in-training) 227 (-52) 110 (-14) 72 (-11) 61 (-5) 68 (-14)
Table IV presents the minimum bandwidth overhead achieved by each defense method under varying fixed time overhead thresholds. Adaptive Tamaraw on in-training webpages shows substantial gains over the original Tamaraw, especially under tight time constraints. For instance, when the time overhead is limited to 10%, Adaptive Tamaraw (in-training) achieves a 52% reduction in bandwidth overhead compared to the baseline. Notably, these improvements are more pronounced at lower time overhead thresholds, which is particularly practical: minimizing additional delay is often crucial for preserving user experience.
In the out-of-training setting, where all test traces are from previously unseen webpages, Adaptive Tamaraw still manages to slightly outperform the global Tamaraw strategy. The reductions are modest (e.g., 2–7%), but consistent across all thresholds. This suggests that even for webpages not present during training, partial trace similarity to known anonymity sets can still yield meaningful gains.
In addition to the minimum bandwidth values, we report the average bandwidth overhead, computed as the area under the curve (AUC) across all Pareto pairs, normalized by the range of time overheads. Both in-training and out-of-training variants of Adaptive Tamaraw demonstrate reduced average bandwidth overhead compared to the original Tamaraw baseline.
This experiment demonstrates that when Adaptive Tamaraw is trained on webpages the user would visit, it achieves its best performance, but it still outperforms Tamaraw even if it is not.
VI-D Attacker Accuracy
VI-D1 Theoretical Bounds
In this section, we provide a numerical evaluation of the average success probability bound stated in Theorem C.2, which measures the level of security offered by Adaptive Tamaraw against any passive website fingerprinting adversary. We use the anonymity sets generated from the in-training evaluation (Section VI-B).
For each fixed combination of anonymity set size $k$ and Tamaraw’s bucket length parameter $L$, we consider the corresponding anonymity sets and evaluate the success probability bound for all 33 Pareto-optimal $(\rho_{\text{in}},\rho_{\text{out}})$ pairs identified in Section VI-A. For each pair, we compute the right-hand side of the inequality in Theorem V.1. We then average these bounds across all 33 configurations to obtain a success bound for each $(k,L)$ pair.
Figure 6 presents the computed theoretical upper bounds on the attack success probability, as described in Theorem C.2, across varying anonymity set sizes $k$ and for three different values of the bucket length parameter $L\in\{100,500,1000\}$.
As expected, the attacker’s success probability decreases consistently with larger anonymity set sizes. This confirms that increasing $k$ enhances the uncertainty from the adversary’s perspective, thereby reducing the effectiveness of any passive WF attack. It is interesting to note that the decrease is not inverse to $k$; which says that larger $k$ encourages anonymity sets that have many traces from the same site. Additionally, increasing $L$ (padding traces to longer lengths) also suppresses maximum attacker accuracy. For instance, at $k=7$, the success probability drops from approximately 45% with $L=100$ to about 35% with $L=1000$, but note that the overhead cost is substantial as seen earlier in Table III.
Figure 6: Formal upper bound on an attacker’s accuracy for different anonymity-set sizes $k$ and bucket-length parameters $L$ in our Adaptive Tamaraw analysis (Section V-F). Increasing $k$ enlarges the pre-image of each defended trace, while a larger $L$ suppresses size-based leakage within each bucket. Together these levers tighten the bound from $\approx\!0.75$ at $(k,L)=(2,100)$ to below $0.17$ at $(k,L)=(30,1000)$, illustrating the trade-off between security and overhead.
We also evaluated the theoretical bound in the out-of-training scenario with $k=7$ and $L=100$. Specifically, we assigned each trace from the out-of-training dataset to its predicted anonymity set and computed the corresponding theoretical bound, then averaged the results over all traces. This yielded a maximum attacker accuracy bound of 31%, indicating that the theoretical framework remains valid even when test traces originate from webpages unseen during training. Notably, this bound is lower than its in-training counterpart in Figure 6 (45%), which can be attributed to the defense behavior in the out-of-training setting: as shown by the modest bandwidth improvements in Table IV, the defense tends to fall back toward behavior resembling the original, non-adaptive Tamaraw. As a result, the applied parameters are more uniform across traces, leading to more homogeneous traffic patterns and lower attack success rates.
VI-D2 Empirical Attacks
To validate the theoretical success probability bounds derived in the previous section, we evaluate the effectiveness of state-of-the-art WF attacks on traces defended by Adaptive Tamaraw, and compare against the corresponding theoretical upper bounds to assess the tightness of the bound presented in Theorem 6.2. We consider four widely-used and high-performing WF attack models: kFP [4], Tik-Tok  [7], RF  [8], and LASERBEAK [9].
In alignment with our deployment-oriented evaluation in Section VI-C, we use the practical configuration of $k=7$ and $L=100$. We randomly sample four $(\rho_{\text{in}},\rho_{\text{out}})$ pairs from the 33 Pareto-optimal global configurations and apply Adaptive Tamaraw to the training traces using each selected pair. We then conduct adversarial training using the defended traces and evaluate the attack success rate of each WF model on the corresponding test set. The training protocols and hyperparameters for all four models are the same mentioned in their original paper.
Table V compares the empirical classification accuracy of the four WF attack models against the theoretical upper bounds. Across all evaluated cases, we observe that the empirical success rate of each model remains consistently below the theoretical bound, where RF is the model achieving the closest results to the bound. This shows that the bound provided in Theorem 6.2 is useful as a principled estimate of attack performance.
TABLE V: Comparison of the theoretical maximum attacker accuracy bound with the empirical accuracy of various website fingerprinting attacks (kFP, Tik-Tok, RF, LASERBEAK) across four randomly selected $(\rho_{\text{out}},\rho_{\text{in}})$ padding rate pairs. Results are shown for $k=7$ and $L=100$.
Padding rates Accuracy (%)
$\rho_{\text{out}}$ $\rho_{\text{in}}$ Bound kFP Tik-Tok RF LASERBEAK
0.012 0.030 41 31 20 39 31
0.090 0.010 43 33 21 38 32
0.010 0.020 42 31 22 40 30
0.030 0.120 35 20 15 33 23
VII Discussion and Conclusion
We introduced a new defense framework that bridges regularization and supersequence methods with three core components: pattern-level clustering, $(k,l)$-diverse anonymity set construction, and early-time switching. We instantiated this design using Tamaraw, yielding Adaptive Tamaraw, a defense that maintains provable guarantees while adapting to trace-specific structure.
Based on our experiments, adaptive Tamaraw offers meaningful gains when tested on webpages it was trained on, and performs similarly to the original Tamaraw on webpages it was not trained on. As $k$ increases, the defense gradually resembles the classic Tamaraw, allowing control over the trade-off between privacy and overhead (Table  III and Figure 6).
While we instantiate our approach using Tamaraw due to its formal upper bound on information leakage, the core our approach is general and can be extended to other defenses that support fixed reugalarization configurations.
Appendix A Model Architectures
Holmes. We adopt the Holmes model architecture proposed in [40] for early-stage website fingerprinting. Holmes utilizes a hybrid convolutional encoder trained via supervised contrastive learning to align partial traffic segments with their corresponding full-trace embeddings. The encoder consists of two initial 2D convolutional blocks, each composed of a Conv2D layer, batch normalization, and max pooling, followed by four 1D convolutional blocks comprising Conv1D layers with kernel size 3, batch normalization, and pooling operations. Residual connections are incorporated throughout the network to preserve spatial-temporal dependencies during downsampling. The model learns to embed partial traces such that their representations closely match those of complete traces from the same website, to achieve early prediction with limited input.
[…]
Appendix B Additional Experimental Results
B-A Tamaraw Parameter Grid and Pareto Frontier
To characterize the trade-off between bandwidth and time overhead in the Tamaraw defense, we constructed a grid of $(\rho_{\text{in}},\rho_{\text{out}})$ pairs following the method recommended in the original Tamaraw paper [11]. Specifically, we started with the original values used in that work:
$\rho_{\text{in}}^{\text{init}}=0.012,\quad\rho_{\text{out}}^{\text{init}}=0.04,$
and varied each parameter independently across 14 values, ranging from 7 times smaller to 7 times larger on a logarithmic scale. That is, each consecutive value differs from the previous by a multiplicative factor of $e^{\ln(7)/7}\approx 1.316$. This results in a total of $14\times 14=196$ combinations of $(\rho_{\text{in}},\rho_{\text{out}})$ pairs.
For each configuration, we applied the Tamaraw defense and measured both the bandwidth and time overhead using a batch of traces, fixing the parameter $L=100$. Among the resulting overhead pairs, we retained only the Pareto-optimal configurations (those for which no other parameter pair simultaneously achieves lower bandwidth and time overhead). The 33 obtained points form the Pareto frontier, representing the optimal trade-off curve between efficiency and latency.
Figure 7 illustrates the resulting Pareto frontier. Each red dot corresponds to a Pareto-optimal $(\rho_{\text{in}},\rho_{\text{out}})$ configuration, annotated with its corresponding (bandwidth, time) overhead pair. The curve reveals the inherent trade-off: significant reductions in one metric typically come at the cost of increases in the other.
Figure 7: Pareto frontier for Tamaraw overheads computed from 196 parameter pairs for $L=100$. Each point corresponds to a $(\rho_{\text{in}},\rho_{\text{out}})$ configuration that is Pareto-optimal with respect to bandwidth and time overhead.
B-B Overhead Improvement Across Cluster Sizes
To evaluate the effectiveness of adaptive switching, we measure the overhead reduction achieved by Adaptive Tamaraw compared to the static global Tamaraw defense. For each test trace, we first compute the total overhead (sum of bandwidth and time overhead) under both the global configuration and the adaptive configuration, and report the average difference across all traces. Figure 8 illustrates the average overhead improvement achieved by Adaptive Tamaraw over static global Tamaraw configurations, plotted as a function of the anonymity set size $k$. Results are shown for three different bucket-length parameters: $L\in\{100,500,1000\}$. The vertical axis reports the average total overhead reduction across all test traces and global configurations. Across all values of $L$, we observe that Adaptive Tamaraw consistently reduces overhead compared to the baseline. Smaller anonymity sets (i.e., decreasing $k$) generally lead to greater improvements, especially when $L$ is large. For $L=100$, the average overhead improvement steadily rises with $k$, reaching a reduction of approximately 15% at $k=30$. A similar but slightly more subdued trend is seen for $L=500$, where the maximum gain is around 25%. The most dramatic improvements are observed for $L=1000$, where average overhead savings exceed 60% at $k=30$ and remain significantly better than the baseline even at higher values of $k$.
This trend highlights the benefit of clustering traffic patterns into larger and more diverse anonymity sets, which allows the second phase of the defense to apply more relaxed parameters while preserving provable security. It also confirms that the adaptive switching strategy is particularly effective when the regularization budget (controlled by $L$) is high—leading to large opportunities for savings through early switching. Interestingly, the improvements are often most pronounced when $k$ is small (reaching around 100% reduction when $L=1000$, and 40% when $L=100$). This may be due to the fact that smaller anonymity sets contain more homogeneous traffic patterns, allowing the anonymity set–specific Tamaraw parameters to better optimize a narrower distribution of traces. As $k$ increases, the diversity within each set grows, which can dilute the effectiveness of a single shared configuration across that set.
Figure 8: Average total overhead improvement of Adaptive Tamaraw compared to global Tamaraw, plotted across different anonymity set sizes $k$ for three values of $L$.
B-C Cluster Diversity Evaluation
As discussed in Section V-D, our proposed $k$-anonymity-based clustering algorithm relies on a Tamaraw-specific distance function that implicitly encourages low intra-cluster diversity. To empirically validate this behavior, we evaluate the purity of the anonymity sets generated by our clustering algorithm.
Following standard definitions [43], we compute the purity $p_{i}$ of the $i$-th cluster as:
$p_{i}\triangleq\frac{\max_{j}N_{i,j}}{N_{p_{i}}}\times 100\%,$
where $N_{i,j}$ is the number of traces in cluster $i$ that originate from website $j$, and $N_{p_{i}}$ is the total number of traces in cluster $i$. The purity thus reflects the dominance of the most frequent website in a given cluster, with lower values indicating greater diversity.
We compute the average purity across all clusters for each anonymity set size $k$ with fixed $L=100$. Figure 9 presents the resulting values alongside the theoretical lower bound $f(k)=1/k$, which represents the best achievable diversity for a perfectly balanced cluster.
Figure 9: Average cluster purities vs. anonymity set size $k$, for $L=100$. The dashed line shows the theoretical lower bound $1/k$.
As shown in the figure, the empirical purity values closely track the $1/k$ curve, indicating that our clustering procedure approaches the theoretical optimum in terms of diversity. This result confirms that the Tamaraw-specific distance function naturally leads to anonymity sets composed of traffic patterns from multiple websites, supporting the desired $l$-diversity property of the defense.
Appendix C Proofs
C-A Security Bound for Adaptive Tamaraw
Notation.  Let $\mathcal{S}=\{S_{1},\dots,S_{m}\}$ be the anonymity sets of traffic patterns (produced in Section V-D). For every set $S_{i}$ we fix a safe time $t_{S_{i}}$; this value is the same for all traces in $S_{i}$ but may differ across sets.
We use two padding operators:
  • $D_{\text{glob}}(\,\cdot\,;\,p_{\mathrm{in}}^{g},p_{\mathrm{out}}^{g})$ applies the global constant–rate schedule with parameters $p_{\mathrm{in}}^{g},p_{\mathrm{out}}^{g}$.
  • $D_{S_{i}}(\,\cdot\,;\,p_{\mathrm{in}}^{(S_{i})},p_{\mathrm{out}}^{(S_{i})},L)$ applies the AS–specific schedule chosen for $S_{i}$.
Given a raw trace $f\in S_{i}$ the full defense acts in two stages:
$\widehat{f}=D(f)=\begin{cases}D_{\text{glob}}\!\bigl{(}f(t);\;p_{\mathrm{in}}^{g},p_{\mathrm{out}}^{g}\bigr{)},&0\leq t<t_{S_{i}},\\[6.0pt] D_{S_{i}}\!\bigl{(}f(t);\;p_{\mathrm{in}}^{(S_{i})},p_{\mathrm{out}}^{(S_{i})},L\bigr{)},&t\geq t_{S_{i}}.\end{cases}$
We write this concisely as a labelled transition
$D_{\text{glob}}\;\xrightarrow[\;t_{S_{i}}\;]{}D_{S_{i}}.$
Because $t_{S_{i}}$ is fixed inside the set, an observer who notices the switch time learns only that the trace belongs to an anonymity set whose safe time equals $t_{S_{i}}$.
AS-specific tail. Inside an anonymity set $S_{i}$, let
$S_{i,\ell}=\left\{t\in S_{i}:\operatorname{len}(t)=\ell\right\}$
denote the subset of defended traces whose total length is $\ell$. Each $S_{i,\ell}$ corresponds to a bucket of traces that become indistinguishable after regularization. For each such bucket, an optimal attacker performs a majority vote, achieving success rate
$\frac{\max_{w}\left|\{t\in S_{i,\ell}:\operatorname{site}(t)=w\}\right|}{|S_{i,\ell}|}.$
To model the average-case attacker success over $S_{i}$, we define the non-uniform attacker accuracy as the expected success rate across all observed length buckets. Let $\mathcal{L}$ denote the set of all distinct length buckets observed in $S_{i}$. Then:
$\bar{A}(S_{i})$ $=\sum_{\ell\in\mathcal{L}}\frac{|S_{i,\ell}|}{|S_{i}|}\cdot\frac{\max\limits_{w}\left|\left\{t\in S_{i,\ell}:\operatorname{site}(t)=w\right\}\right|}{|S_{i,\ell}|}$ (2)
$=\mathbb{E}_{\ell\sim\mathcal{L}}\left[\frac{\max\limits_{w}\left|\left\{t\in S_{i,\ell}:\operatorname{site}(t)=w\right\}\right|}{|S_{i,\ell}|}\right]$ (3)
Lemma C.1 (Post-switch non-uniform weighted $\delta$).
For every anonymity set $S_{i}$, the AS-specific mapping $D_{S_{i}}$ is non-uniformly $\delta_{i}$-weighted-non-injective, where $\delta_{i}={1}/{\bar{A}(S_{i})}$ and $\bar{A}(S_{i})$ is the average attacker success rate defined in (2).
Proof.
After the switch, all traces in $S_{i}$ are padded using the fixed parameters $(p_{\mathrm{in}}^{(S_{i})},p_{\mathrm{out}}^{(S_{i})},L)$, so the only observable feature is the total defended length $\ell$. This partitions $S_{i}$ into buckets
$S_{i,\ell}=\left\{t\in S_{i}:\operatorname{len}(t)=\ell\right\}.$
Weighted pre-image size. Each bucket length $\ell$ (i.e., each trace $t\in S_{i,\ell}$) corresponds to a single defended trace $f^{\prime}_{i,\ell}$ . The weighted pre-image size of $f^{\prime}_{i,\ell}$, by (1), is
$\tilde{\delta}\left(f^{\prime}_{i,\ell}\right)=\frac{|S_{i,\ell}|}{\displaystyle\max_{w}\left|\left\{t\in S_{i,\ell}:\text{site}(t)=w\right\}\right|}.$
Attacker success. The optimal attacker outputs the majority website in each bucket, achieving success $1/\tilde{\delta}(f^{\prime}_{i,\ell})$.
Average-case accuracy. The attacker’s expected success rate over $S_{i}$ is given by the non-uniform accuracy
$\bar{A}(S_{i})=\sum_{\ell}\frac{|S_{i,\ell}|}{|S_{i}|}\cdot\frac{1}{\tilde{\delta}(f^{\prime}_{i,\ell})}=\mathbb{E}_{f^{\prime}\sim S_{i}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right],$
Therefore, by the definition in Section IV, the mapping $D_{S_{i}}$ is non-uniformly $\delta_{i}$-weighted-non-injective with $\delta_{i}=1/\bar{A}(S_{i})$. It follows that the attacker’s overall success rate for any trace in $S_{i}$ is bounded by
$\Pr[\text{success}\mid S_{i}]\;\leq\;\bar{A}(S_{i}).\qed$
Theorem C.2 (Global non-uniformly weighted $\delta$–non-injectivity).
Let $\mathcal{S}$ be the set of all anonymity sets obtained, and for each anonymity set $S_{i}\in\mathcal{S}$, let
$\delta_{i}=\frac{1}{\bar{A}(S_{i})}.$
Define the global non-uniform weighted injectivity parameter $\delta$ as
$\frac{1}{\delta}=\mathbb{E}_{S_{i}\sim\mathcal{S}}\left[\frac{1}{\delta_{i}}\right]=\mathbb{E}_{S_{i}}[\bar{A}(S_{i})].$
Then, Adaptive Tamaraw is non-uniformly weighted $\delta$-non-injective, and the attacker’s average success probability is bounded by
$\boxed{\Pr[\text{success}]\;\leq\;\frac{1}{\delta}}.$
Proof.
Recall from Lemma C.1 that for each anonymity set $S_{i}$, the attacker’s success rate over its defended outputs is bounded by $\bar{A}(S_{i})=\mathbb{E}_{f^{\prime}\sim S_{i}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right]$, and thus the mapping $D_{S_{i}}$ is non-uniformly $\delta_{i}$-weighted-non-injective, where $\delta_{i}=1/\bar{A}(S_{i})$.
Now consider the overall defended distribution $\mathcal{F}^{\prime}$, which consists of all post-regularization traces $f^{\prime}_{i,\ell}$ produced from inputs in various anonymity sets $S_{i}$. Each defended trace $f^{\prime}$ must originate from some $S_{i}$ with probability $P(S_{i})$, and within $S_{i}$, from a bucket of length $\ell$ with probability $|S_{i,\ell}|/|S_{i}|$.
We begin with the average inverse diversity over anonymity sets:
$\mathbb{E}_{S_{i}}\left[\frac{1}{\delta_{i}}\right]$ $=\mathbb{E}_{S_{i}}\left[\bar{A}(S_{i})\right]$
$=\sum_{i}P(S_{i})\cdot\bar{A}(S_{i})$
$=\sum_{i}P(S_{i})\cdot\sum_{\ell}\frac{|S_{i,\ell}|}{|S_{i}|}\cdot\frac{1}{\tilde{\delta}(f^{\prime}_{i,\ell})}.$
Note that $P(S_{i})\cdot\frac{|S_{i,\ell}|}{|S_{i}|}$ is precisely the probability of observing the defended trace $f^{\prime}_{i,\ell}$ in the output space $\mathcal{F}^{\prime}$, since:
$P(f^{\prime}_{i,\ell})=P(S_{i})\cdot\frac{|S_{i,\ell}|}{|S_{i}|}.$
Also, $\tilde{\delta}(f^{\prime}_{i,\ell})$ is the weighted pre-image size of $f^{\prime}_{i,\ell}$ as defined in Section IV.
We can therefore rewrite the sum as:
$\sum_{i}\sum_{\ell}P(S_{i})\cdot\frac{|S_{i,\ell}|}{|S_{i}|}\cdot\frac{1}{\tilde{\delta}(f^{\prime}_{i,\ell})}$ $=\sum_{f^{\prime}\in\mathcal{F}^{\prime}}P(f^{\prime})\cdot\frac{1}{\tilde{\delta}(f^{\prime})}$
$=\mathbb{E}_{f^{\prime}\sim\mathcal{F}^{\prime}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right].$
Thus, we have:
$\mathbb{E}_{S_{i}}\left[\frac{1}{\delta_{i}}\right]=\mathbb{E}_{f^{\prime}\sim\mathcal{F}^{\prime}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right],$
which confirms that the expected attacker success over anonymity sets equals the expected attacker success over the defended output distribution.
Therefore, Adaptive Tamaraw as a whole satisfies the non-uniform $\delta$-non-injectivity condition with
$\delta=\left(\mathbb{E}_{f^{\prime}\sim\mathcal{F}^{\prime}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right]\right)^{-1},$
completing the proof. ∎
Appendix D Modified CAST Algorithm
This section presents the algorithmic components used in Adaptive Tamaraw. In particular, we provide the full pseudocode for our modified version of the CAST algorithm, which plays a central role in obtaining the intra-website patterns, mentioned in Section V-C.
Algorithm 2 Intra-Website Pattern Detection via Modified CAST
1:A set of traffic traces $\{F_{1},F_{2},\ldots,F_{n}\}$ from a single website
2:A set of balanced clusters $\mathcal{P}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{m}\}$
3:Step 1: Similarity Matrix Construction
4:for each trace $F_{i}$ do
5:  Compute Euclidean distances $d(F_{i},F_{j})$ for all $j$
6:  Determine local scale $\sigma_{i}$ as the distance from $F_{i}$ to its $K$th nearest neighbor (with $K=7$)
7:end for
8:Compute the similarity matrix $S$, where:
[…]
10:Step 2: Initial Clustering with Modified CAST
11:Initialize all traces as unassigned.
12:while there exist unassigned traces do
13:  for each unassigned trace $F_{x}$ do
14:   for each existing cluster $\mathcal{C}$ do
15:     Compute affinity:
16:     $a_{\mathcal{C}}(F_{x})=\frac{1}{|\mathcal{C}|}\sum_{F_{y}\in\mathcal{C}}S(F_{x},F_{y})$
17:   end for
18:   Assign $F_{x}$ to the cluster $\mathcal{C}^{*}$ for which $a_{\mathcal{C}^{*}}(F_{x})$ is maximum, provided that
19:   $a_{\mathcal{C}^{*}}(F_{x})\geq\tau$, where $\tau$ is a dynamically chosen affinity threshold.
20:   if no existing cluster meets the threshold then
21:     Create a new cluster with $F_{x}$ as the seed.
22:   end if
23:  end for
24:  Perform a cleaning step: For every trace, reassign it to the cluster with which it has the highest affinity if different from its current cluster.
25:end while
26:Step 3: Post-Processing
27:while the total number of clusters $\text{MaxClusters}$ do
28:  Sort clusters in $\mathcal{P}$ by size.
29:  Merge the smallest cluster into the adjacent cluster that minimizes the expansion ratio, defined as
30:  $\phi(\mathcal{C})=\frac{\mathrm{cut}(\mathcal{C})}{\mathrm{vol}(\mathcal{C})}$,
31:  where for a cluster $\mathcal{C}$,
32:  $\mathrm{cut}(\mathcal{C})=\sum_{F\in\mathcal{C}}\sum_{F^{\prime}\notin\mathcal{C}}d(F,F^{\prime})$,
33:  and
34:  $\mathrm{vol}(\mathcal{C})=\sum_{F,F^{\prime}\in\mathcal{C}}d(F,F^{\prime})$.
35:  Update the clustering accordingly.
36:end while
37:return the final set of clusters $\mathcal{P}$.